OpenStudy (anonymous):

log of 3squareroot of 5 base 5???

OpenStudy (anonymous):

\[\log_{5} 3\sqrt{5}\]?

OpenStudy (anonymous):

=1.1826061944859852951345663592710522530246693998731672096600...

OpenStudy (anonymous):

\[\log_{5} 3\sqrt{5} = x \] apply change of base formula, recall... \[\log_{A} B = (\log_{10} B )\div(\log_{10}A) \] a funny analogy is that B is the bigger one right? so B is the BIG WIFE A is the base, visually smaller than B the big wife right? :) A is the concubine, So when you apply change of base, the big wife will forever be upstairs (bigger) and the concubine will always be below, silly but i hope its funny and helps to remember. therefore \[\log_{5} (3\sqrt{5}) = (\log_{10}(3\sqrt{5})) \div (\log_{10}(5)) \] You change it to base 10 because your calculator can only calculate when it is log base 10 or log base e (ln)

OpenStudy (anonymous):

it would be useful to convert into exponential form, because it is hard to evaluate this :)